I rcvd an answer to this problem earlier (24882) but the steps and the reason why are still not clear to me. This was not taught to us as a formula, but just as a regular equation. So I am not sure why you would reverse the fraction instead of multiplying everything by 5 to remove the fraction
HELP!!!
F=9/5c+32, solve for C

Answer by AnlytcPhil(1806)

I rcvd an answer to this problem earlier (24882) but the steps and the reason
why are still not clear to me. This was not taught to us as a formula, but just
as a regular equation. So I am not sure why you would reverse the fraction
instead of multiplying everything by 5 to remove the fraction
HELP!!! 
F=9/5c+32, solve for C

You can multiply everything by 5

F = 9/5C + 32

5F = 9C + 160

5F - 160 = 9C

Factor out 5 on the left

5(F - 32) = 9C

Divide both sides by 9

5/9(F - 32) = C

or

C = 5/9(F - 32)

Edwin
AnlytcPhil@aol.com